OpenStudy (anonymous):

Solve for p. 3(p+q)=p (1 point) A. q=-2/3p B. q=-3/2p C. p=2/3q D. p=3/2q

OpenStudy (anonymous):

3(p+q)=p expanding... 3p + 3q = p collecting like terms... 3q = p - 3p = -2p -2p = 3q div both sides by -2 p = -3/2q